OpenStudy (anonymous):

Find the roots of the polynomial equation: x2+2x-3=0

OpenStudy (anonymous):

X=-3 x=1

OpenStudy (anonymous):

x (x+2) = 3

OpenStudy (anonymous):

(x+3)(x-1)=0 x+3 = 0 x = -3 x-1= 0 x= 1 so x = 1 or -3